Find the exact value of cot(sin-1 7/9)
I really have no idea, I was looking for a value of 7/9 on my table but I can't find it.

Answer by scott8148(6628)

use a right triangle for a template

for angle A (the angle whose sin is 7/9) ; the side opposite the angle is 7 and the hypotenuse is 9

using Pythagoras, find the other side (adjacent to A) ___ x^2 = 9^2 - 7^2 = 32 ___ x = 4√2

cot(A) = adjacent / opposite = (4√2) / 7